[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Aug 16, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Aug 17 02:32:33 UTC 2017


HSU Shift Log for Wed Aug 16 2017 16:29:12 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues: Microwave images are not displaying on the ATCF for Gert, but  
works fine for AL91 and AL92. Was able to do it at the beginning of the  
shift but it stopped working sometime during the morning. Closing the  
window did not help, and it's doing it on all the workstations.

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170817/1cb19564/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list